OpenStudy (anonymous):

square root of 32x^10 divided by square root 2x^4

OpenStudy (anonymous):

\[\sqrt {32x^{10}}/\sqrt{2x^4}\]The square roots cancel each other, so we are left with \[32x^{10}/2x^4\]\[16x^{10}/x^4\]\[16x^6\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[\frac{\sqrt {32x^{10}}}{\sqrt{2x^4}}\] \[\sqrt{\frac{32x^{10}}{2x^4}}=\sqrt{16x^6}=4x^3\]